one number is 32 more than the second number twice the first number is 4 more than 8 times the second number. Find the numbers.
Answer by lwsshak3(11628)

one number is 32 more than the second number twice the first number is 4 more than 8 times the second number. Find the numbers.
**
let x= second number
x+32=first number
..
2(x+32)=8x+4
2x+64=8x+4
6x=60
x=10
x+32=42
second number=10
first number=42